OpenStudy (anonymous):

The sum of three consecutive integers is 267. What are the three integers?

OpenStudy (anonymous):

let N be your smallest integer N + (N+1) + (N+2) = 267

OpenStudy (anonymous):

3N + 3 = 267 3N = 264 N = 88 CHECK: 88 + 89 + 90 = 267

OpenStudy (anonymous):

\[x+(x+1)+(x+2)=267\]\[3x+3=267\] \[3x= 264\] \[x= 88\] numbers are 88, 89, 90